\begin{thebibliography}{99}

\bibitem{Marshall}
A.W. Marshall and I. Olkin.
\newblock Inequalities: Theory of Majorization and Its Application.
\newblock Academic Press, New York (1979).

\bibitem{Nielsen}
M. Nielsen and Guifr\'{e} Vidal.
\newblock Majorization and the interconversion of bipartite states.
\newblock Quant. Inf. \& Comput. \textbf{1}(1), 76-93(2001).

\bibitem{Hiroshima}
T. Hiroshima.
\newblock Majorization criterion for distillability of a bipartite quantum state.
\newblock \href{http://dx.doi.org/10.1103/PhysRevLett.91.057902}{Phys. Rev. Lett. \textbf{91}, 057902 (2003).}

\bibitem{Partovi}
M.H. Partovi.
\newblock Entanglement detection using majorization uncertainty bounds.
\newblock \href{http://dx.doi.org/10.1103/PhysRevA.86.022309}{Phys. Rev. A \textbf{86}, 022309 (2012).}

\bibitem{Watrous} J.~Watrous, Theory of Quantum Information,
University of Waterloo, Waterloo (2008). See
\url{http://www.cs.uwaterloo.ca/~watrous/quant-info/}


\bibitem{Hardy}
G.H. Hardy, J.E. Littlewood and G. Polya.
\newblock Inequalities (2ed)
\newblock Cambridge University Press (1932).

\bibitem{Ando}
T. Ando.
\newblock Doubly stochastic matrices, and comparison of eigenvalues.
\newblock \href{http://dx.doi.org/10.1016/0024-3795(89)90580-6}{Linear Algebra Appl. \textbf{118}, 163-248 (1989).}

\bibitem{Hiai}
F. Hiai.
\newblock Similarity preserving linear maps on matrices.
\newblock \href{http://dx.doi.org/10.1016/0024-3795(87)90145-5}{Linear Algebra Appl. \textbf{97}, 127-139 (1987).}

\bibitem{Lindblad}
G.~Lindblda.
\newblock Completely positive maps and entropy inequalities.
\newblock \href{http://dx.doi.org/10.1007/BF01609396}{Commun. Math. Phys. \textbf{40}, 147-151 (1975).}

\bibitem{Li}
Y. Li and Y. Wang.
\newblock Further results on entropy and separability.
\newblock \href{http://dx.doi.org/10.1088/1751-8113/45/38/385305}{J. Phys. A: Math. Theor. \textbf{45}, 385305 (2012).}

\bibitem{Zhang}
L. Zhang and J. Wu.
\newblock Von Neumann entropy-preserving quantum operations.
\newblock \href{http://dx.doi.org/10.1016/j.physleta.2011.10.008}{Phys. Lett. \textbf{375}(47): 4163-4165 (2011).}

\end{thebibliography}

\documentclass[12pt]{article}

\usepackage[T1]{fontenc}
\usepackage[sc]{mathpazo}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{enumerate}
\usepackage{amsthm}% theorems and proofs
\usepackage{amsfonts,mathrsfs}
\usepackage[style]{fncychap}
\usepackage{graphicx} % For including graphics N.B. pdftex graphics driver
\usepackage{geometry} % allows easy specifying of the page layout
\usepackage{eepic}
\usepackage{ifthen}
\newboolean{ElectronicVersion}
\setboolean{ElectronicVersion}{true} % CHANGE THIS AS REQUIRED

\usepackage{hyperref}
\hypersetup{pdfpagemode=UseNone}

\newcommand{\op}[1]{\operatorname{#1}}

\renewcommand{\t}{{\scriptscriptstyle\mathsf{T}}}

\newcommand{\setft}[1]{\mathrm{#1}}
\newcommand{\lin}[1]{\setft{L}\left(#1\right)}
\newcommand{\density}[1]{\setft{D}\left(#1\right)}
\newcommand{\unitary}[1]{\setft{U}\left(#1\right)}
\newcommand{\trans}[1]{\setft{T}\left(#1\right)}

\newcommand{\rank}{\op{rank}}

\newcommand{\Sym}{\operatorname{Sym}}

\newcommand{\locc}{\operatorname{LOCC}}

\def\complex{\mathbb{C}}
\def\real{\mathbb{R}}

\def\I{\mathbb{1}}

\newenvironment{mylist}[1]{\begin{list}{}{
    \setlength{\leftmargin}{#1}
    \setlength{\rightmargin}{0mm}
    \setlength{\labelsep}{2mm}
    \setlength{\labelwidth}{8mm}
    \setlength{\itemsep}{0mm}}}
    {\end{list}}

\def\ot{\otimes}

\newcommand{\out}[2]{| #1\rangle\langle #2 |}

\newcommand{\defeq}{\stackrel{\smash{\textnormal{\tiny def}}}{=}}

\newcommand{\Herm}{\mathrm{Herm}}

\newcommand{\pa}[1]{(#1)}
\newcommand{\Pa}[1]{\left(#1\right)}

\newcommand{\Br}[1]{\left[#1\right]}
\newcommand{\set}[1]{\{#1\}}

\newcommand{\ket}[1]{|#1\rangle}

\def\Jamiolkowski{J}
\newcommand{\jam}[1]{\Jamiolkowski\pa{#1}}

\DeclareMathOperator{\vectorize}{vec}
\newcommand{\col}[1]{\vectorize\pa{#1}}
\newcommand{\row}[1]{\vectorize\pa{#1}^{\dagger}}

\DeclareMathOperator{\trace}{Tr}

\newcommand{\Ptr}[2]{\trace_{#1}\Pa{#2}}

\newcommand{\Tr}[1]{\Ptr{}{#1}}

\def\cH{\mathcal{H}}

\def\rS{\mathrm{S}}\def\rT{\mathrm{T}}

\newtheorem{thrm}{Theorem}[section]

\newtheorem{prop}[thrm]{Proposition}

\theoremstyle{definition}

\numberwithin{equation}{section}

\newcounter{questionnumber}

\begin{document}

%========================================================================================%
\title{Majorization-preserving quantum channels}
%========================================================================================%

\author{Lin Zhang\footnote{E-mail: godyalin@163.com;
linyz@hdu.edu.cn}\\
  {\it\small Institute of Mathematics, Hangzhou Dianzi University, Hangzhou 310018, PR~China}}

\date{}
\maketitle \mbox{}\hrule\mbox\\
\begin{abstract}

In this report, we give a characterization to those quantum channels
that preserve majorization relationship between quantum states. Some remarks are presented as well.

\end{abstract}
\mbox{}\hrule\mbox\\

%=============================================================================%
\section{Introduction}
%=============================================================================%

The notion of majorization has been widely used in many fields of
mathematics, such as matrix analysis, operator theory, frame theory,
and inequalities involving convex functions. The readers can be
referred to a standard literature by Marshall and Olkin
\cite{Marshall}.

This report discusses the notion of majorization and some of its
connections to quantum information. As a tool used in quantum
information theory, the main application of majorization is first
given by M. Nielsen \cite{Nielsen}. His result precisely
characterizes when it is possible for two parties to transform one
pure state into another by means of local operations and classical
communication ($\locc$). There are other interesting applications of
the notion can be found in \cite{Hiroshima,Partovi}.

%=============================================================================%
\section{Preliminaries}
%=============================================================================%

Let $\cH$ be a finite dimensional complex Hilbert space. A
\emph{quantum state} $\rho$ on $\cH$ is a positive semi-definite
operator of trace one, in particular, for each unit vector
$\ket{\psi} \in \cH$, the operator $\rho = \out{\psi}{\psi}$ is said
to be a \emph{pure state}. Now the set of all linear operators acting on $\cH$ denoted by $\lin{\cH}$. We denote by $\Herm(\cH)$ the set of all Hermitian operators in $\lin{\cH}$. The notation $\unitary{\cH}$ stands for the set of all unitary operator in $\lin{\cH}$. The set of all quantum states on $\cH$ is
denoted by $\density{\cH}$. For each quantum state
$\rho\in\density{\cH}$, its von Neumann entropy is defined by
$\rS(\rho) = - \Tr{\rho\log_2\rho}$. A \emph{quantum channel}
$\Phi$ on $\cH$ is a trace-preserving completely positive linear
map defined over the set $\density{\cH}$. It follows from
(\cite[Prop.~5.2 and Cor.~5.5]{Watrous}) that there exists linear
operators $\set{K_\mu}_\mu$ on $\cH$ such that $\sum_\mu
K^\dagger_\mu K_\mu = \I$ and $\Phi = \sum_\mu \mathrm{Ad}_{K_\mu}$,
that is, for each quantum state $\rho$, we have the Kraus
representation
\begin{eqnarray*}
\Phi(\rho) = \sum_\mu K_\mu \rho K^\dagger_\mu.
\end{eqnarray*}
The \emph{Choi-Jamio{\l}kowski isomorphism} for a quantum channel $\Phi\in\trans{\cH}$ is defined as
$$
\jam{\Phi} \defeq \Phi\ot\I_{\lin{\cH}}(\col{\I}\row{\I}),
$$
where $\col{\I} \defeq \sum_j\ket{jj}$. In fact, $\jam{\Phi} = \sum_\mu \col{K_\mu}\row{K_\mu}$.

\subsection{Majorization for real vectors}

For two vectors $u,v\in\real^n$, $u$ is \emph{majorized} by $v$,
denoted by $u\prec v$, if
$$
\sum^k_{i=1} u^\downarrow_i \leqslant \sum^k_{i=1} v^\downarrow_i;
k=1,\ldots,n
$$
and
$$
\sum^n_{i=1} u^\downarrow_i = \sum^n_{i=1} v^\downarrow_i,
$$
where $u^\downarrow_1\geqslant \cdots \geqslant u^\downarrow_n$ is
the decreasing rearrangement of all components of the vector $u$. In
\cite{Hardy}, Hardy, Littlewood and Polya obtained the following
characterization:
\begin{prop}[\cite{Hardy}]
Let $u,v\in\real^n$. $u$ is majorized by $v$, i.e. $u\prec v$, if
and only if $u = Bv$ for some $n\times n$ bi-stochastic matrix $B$.
\end{prop}
Recall that a matrix $B=[b_{ij}]$ is called \emph{stochastic matrix}
if
$$
b_{ij}\geqslant0\quad\text{and}\quad \sum^n_{i=1} b_{ij}=1.
$$
Furthermore, if the stochastic matrix $B$ satisfy the identity
$\sum^n_{j=1} b_{ij}=1$ as well, then it is called
\emph{bi-stochastic matrix}. Let us write $\Sym(\Sigma)$, for
instance $\Sigma=\set{1,\ldots,n}$, to denote the set of one-to-one
and onto functions of the form $\pi:\Sigma\to\Sigma$ (or, in other
words, the \emph{permutation} of $\Sigma$). For each
$\pi\in\Sym(\Sigma)$, we define a matrix $P_\pi:\real^n\to\real^n$
as
$$
(P_\pi)_{ij} =
\begin{cases}
1, & \text{if}\ i=\pi(j);\\
0, & \text{oherwise}.
\end{cases}
$$
It is clear that every permutation matrix is bi-stochastic, and that
the set of bi-stochastic matrices is a convex set. The following
famous theorem establishes that the bi-stochastic matrices are, in
fact, given by the convex hull of the permutation matrices \cite{Watrous}.

\begin{prop}[The Birkhoff-von Neumann theorem]
Let $B:\real^n\to\real^n$ be a linear map. Then $B$ is a
bi-stochastic matrix if and only if there exists a probability
vector $p \in\real^{n!}$ such that
$$
B = \sum_{\pi\in\Sym(\Sigma)} p(\pi) P_\pi,
$$
where $\Sigma=\set{1,\ldots,n}$.
\end{prop}

We say that $T:\real^n\to\real^n$ preserves the majorization if $Tx\prec Ty$ whenever $x\prec y$, i.e.,
$$
x\prec y\Longrightarrow Tx\prec Ty.
$$
In \cite{Ando}, Ando characterized the structure of linear map over $\real^n$ that preserves the majorization between two vectors.

\begin{prop}[\cite{Ando}]\label{prop:ando} Let $T:\real^n\to\real^n$ be a linear
map. Then $Tu\prec Tv$, whenever $u\prec v$, if and only if one of
the following conditions hold.
\begin{enumerate}[(i)]
\item $Tu = \Tr{u}a$ for some $a\in\real^n$.
\item $Tu = \alpha Pu + \beta \Tr{u}e$ for some $\alpha,\beta\in\real$
and permutation $P:\real^n\to\real^n$.
\end{enumerate}
\end{prop}
Note that throughout this paper the trace of a vector means: $\Tr{u} = \sum_j u_j$.

\subsection{Majorization for Hermitian operators}

We will now define an analogous notion of majorization for Hermitian
operators. For Hermitian operators $A,B\in\Herm(\cH)$, we say that
$A$ is \emph{majorized} by $B$, denoted by $A\prec B$, if there
exists a \emph{mixed unitary channel} $\Phi\in\trans{\cH}$ such that
$A=\Phi(B)$, i.e.
$$
A= \Phi(B) = \sum_j p_j U_j B U_j^\dagger, \quad (\forall j:
p_j\geqslant0;\sum_j p_j=1).
$$
Inspiration for this definition partly comes from the Birkhoff-von
Neumann theorem, the following theorem gives an alternate
characterization of this relationship that also connects it with
majorization for real vectors.

Denote by $\lambda(X)$ the all eigenvalues of $X\in\Herm(\cH)$ arranged as a vector. In particular, $\lambda^\downarrow(X)$ stands for the modified vector $\lambda(X)$ whose components are arranged in decreasing order.

\begin{prop}
Let $A,B\in\Herm(\cH)$. Then $A\prec B$ if and only if
$\lambda(A)\prec\lambda(B)$.
\end{prop}

Concerning linear maps on matrices, the structure of linear maps that preserve a quantity or property
is a hot topic. In particular, for a quantum channel preserving majorization, it is essential for discussing the von Neumann entropy preservation of quantum channels. Denote by $\trans{\cH}$ the set of all linear super-operators
$$
\Phi : \lin{\cH}\to\lin{\cH}.
$$
Specifically, we say that $\Phi\in\trans{\cH}$ preserves the majorization if $\Phi(A)\prec\Phi(B)$ whenever $A\prec B$ over $\Herm(\cH)$, i.e.,
$$
A\prec B\Longrightarrow \Phi(A)\prec\Phi(B).
$$
A similar result to Proposition~\ref{prop:ando} was obtained by Hiai in \cite{Hiai}. This result can be described as follows:

\begin{prop}[\cite{Hiai}]\label{prop:hiai}
Let $\Phi\in\trans{\cH}$ be a linear super-operator. Then the
following statements are equivalent:
\begin{enumerate}[(i)]
\item $\Phi$ preserves majorization, that is, $\Phi(A)\prec
\Phi(B)$ whenever $A\prec B$, where $A,B\in\Herm(\cH)$;
\item either there exists an $A_0\in\Herm(\cH)$ such that
\begin{enumerate}[(a)]
\item $\Phi(X) = \Tr{X}A_0$, for all $X\in\Herm(\cH)$,
\end{enumerate}
or there exist a unitary $U\in\unitary{\cH}$ and
$\alpha,\beta\in\real$ such that $\Phi$ has one of the following
forms:
\begin{enumerate}
\item[(b)] $\Phi(X) = \alpha UXU^\dagger + \beta\Tr{X}\I$, for all
$X\in\Herm(\cH)$;
\item[(c)] $\Phi(X) = \alpha UX^\t U^\dagger + \beta\Tr{X}\I$, for all
$X\in\Herm(\cH)$.
\end{enumerate}
\end{enumerate}
\end{prop}

%=============================================================================%
\section{Main result}
%=============================================================================%

We are interested in the structure of completely positive linear maps (CP) on matrices \cite{Lindblad} since the physical transformations are all represented by CP maps.

Consider the linear super-operator $\Lambda\in\trans{\cH}$ defined as follows:
$$
\Lambda(X) = \frac1n \Tr{X}\I.
$$
This is the \emph{completely depolarizing channel}, which sends every density operator to the maximally mixed state $\frac1n\I$. The Choi-Jamio{\l}kowski representation of $\Lambda$ is
$$
\jam{\Lambda} = \frac1n\I\ot\I.
$$
Now we can state our result:

\begin{thrm}
Let $\Phi\in\trans{\cH}$ be a quantum channel. Then $\Phi(\rho)\prec
\Phi(\sigma)$, whenever $\rho\prec \sigma(\rho,\sigma\in\density{\cH})$, if and only if one of
the following conditions hold.
\begin{enumerate}[(i)]
\item $\Phi(\rho) = \omega$ for some $\omega\in\density{\cH}$.
\item $\Phi(\rho) = \lambda U\rho U^\dagger + (1-\lambda)\frac1n\I$ for some $\lambda\in\Br{-\frac1{n^2-1},1}$
and unitary $U\in\unitary{\cH}$.
\item $\Phi(\rho) = \lambda U\rho^\t U^\dagger + (1-\lambda)\frac1n\I$ for some $\lambda\in\Br{-\frac1{n-1},\frac1{n+1}}$
and unitary $U\in\unitary{\cH}$.
\end{enumerate}
\end{thrm}

\begin{proof}
$(\Longleftarrow)$. This direction of the proof is trivially.\\
$(\Longrightarrow)$. From the Proposition~\ref{prop:hiai}, it follows that, either
$$
\Phi (\rho) = \Tr{\rho} A_0 = A_0.
$$
Since $\Phi$ is a quantum channel, we have that $A_0\geqslant0$ and $\Tr{A_0} =1$. Setting $A_0 =\omega\in\density{\cH}$, we have $\Phi(\rho) = \omega$ for all $\rho\in\density{\cH}$.
Or $\Phi$ must be of the following forms:
$$
\Phi(\rho) = \alpha U\rho U^\dagger + \beta \I\quad\text{or}\quad\Phi(\rho) = \alpha U\rho^\t U^\dagger + \beta \I.
$$
Let
$$
\Psi(\rho) \equiv \alpha\rho + \beta\I\quad\text{or}\quad\Psi(\rho) \equiv \alpha\rho^\t + \beta\I.
$$
Thus $\Phi(\rho) = U\Psi(\rho)U^\dagger$ and $\Psi = \alpha\I_{\lin{\cH}} + \beta\Lambda$ or $\alpha \rT + \beta\Lambda$, where $\rT$ is the transpose super-operator, defined by $\rT(\rho) = \rho^\t$. Therefore $\Phi$ is completely positive if and only if $\Psi$ is completely positive.
By the trace-preservation of $\Psi$, we have that $\alpha + n\beta = 1$. Moreover, we have the Choi-Jamio{\l}kowski representation
$$
\jam{\Psi} = \alpha \col{\I}\row{\I} + \beta \I\ot\I \quad\text{or}\quad \jam{\Psi} = \alpha \jam{\rT} + \beta \I\ot\I.
$$
By the semi-definite positivity of $\jam{\Psi}$, we have
$$
\begin{cases}
n\alpha + \beta &\geqslant0\\
\beta&\geqslant0
\end{cases} \quad\text{or}\quad\begin{cases}
\alpha + \beta &\geqslant0 \\
-\alpha + \beta&\geqslant0
\end{cases}
$$
The above inequalities can be solved via the following identification: $\alpha = \lambda$ and $\beta = \frac{1-\lambda}n$. Since the $\jam{\rT}$ is a self-adjoint and unitary, thus all eigenvalues of $\jam{\rT}$ are $\set{\pm1}$. The desired conclusions can be obtained via some simple calculations.
\end{proof}

%=============================================================================%
\section{Discussion}
%=============================================================================%

Recently, in \cite{Li}, Li obtained the following result:
\begin{prop}
Let $\rho,\sigma\in\density{\cH}$. If $\rho\prec \sigma$ and $\rS(\rho) = \rS(\sigma)$, then $\lambda^\downarrow(\rho) = \lambda^\downarrow(\sigma)$.
\end{prop}
Recall that the von Neumann entropy of a quantum state $\rho$ is defined as $\rS(\rho) = -\Tr{\rho\log_2\rho}$.
This result indicates that, if $\rho\prec \sigma$ and $\rS(\rho) = \rS(\sigma)$, then there exists some unitary $U\in\unitary{\cH}$ such that $\rho = U\sigma U^\dagger$. Furthermore, we can have the following conclusion: if $\rho\prec\sigma$ and $\sigma\prec \rho$, denoted by $\rho\sim\sigma$, then $\rho = U\sigma U^\dagger$ for some unitary $U\in\unitary{\cH}$. Thus if a quantum channel preserves majorization, i.e. $\Phi(\rho)\prec\Phi(\sigma)$ whenever $\rho\prec\sigma$, then we can infer that $\Phi(\rho)\sim\Phi(\sigma)$ whenever $\rho\sim\sigma$. In other words,
$$
\rho = U\sigma U^\dagger \Longrightarrow \Phi(\rho) = V\Phi(\sigma)V^\dagger.
$$
Finally, The characterization of majorization-preserving quantum
channels can be reduced to the characterization of unitary
equivalence preserving quantum channels.

In what follows, we consider the characterization of
majorization-preserving quantum channels for a restricted
majorization class of a fixed quantum state with a complete
different eigenvalues in the qubit case.

\begin{thrm}
Let $\Phi\in\trans{\complex^2}$ be a unital channel and
$\rho_0\in\density{\complex^2}$ be fixed with its both eigenvalues
satisfying that $\lambda_1>\lambda_2>0$. Then $\Phi(\rho)\sim
\Phi(\rho_0)$, whenever $\rho\sim
\rho_0(\rho\in\density{\complex^2})$, if and only if one of the
following conditions hold.
\begin{enumerate}[(i)]
\item $\Phi(\rho) = \frac12\I$.
\item $\Phi(\rho) = \lambda U\rho U^\dagger + (1-\lambda)\frac12\I$ for some $\lambda\in\Br{-\frac13,1}$
and unitary $U\in\unitary{\complex^2}$.
\item $\Phi(\rho) = \lambda U\rho^\t U^\dagger + (1-\lambda)\frac12\I$ for some $\lambda\in\Br{-1,\frac13}$
and unitary $U\in\unitary{\complex^2}$.
\end{enumerate}
\end{thrm}

\begin{proof}
($\Longleftarrow$) This direction of proof is trivial.\\
($\Longrightarrow$) Let $\rho_0 = \lambda_1\out{0}{0} +
\lambda_2\out{1}{1}$. If $\rho\sim \rho_0$, then there exists a
unitary $U$ such that $\rho = U\rho_0U^\dagger$, thus $\rho$ lies in
the unitary orbit of $\rho_0$. By the property of $\Phi$, we see
that
$$
\Phi(U\rho_0U^\dagger) = V\Phi(\rho_0)V^\dagger.
$$
Since $\Phi(\I)=\I$, it follows that
\begin{eqnarray}
\Phi(U\out{0}{0}U^\dagger) &=& V\Phi(\out{0}{0})V^\dagger,\\
\Phi(U\out{1}{1}U^\dagger) &=& V\Phi(\out{1}{1})V^\dagger.
\end{eqnarray}
Thus,
\begin{eqnarray}
\Phi(\out{0}{0})\sim\Phi(\out{u}{u})\sim\Phi(\out{1}{1})
\end{eqnarray}
for any $\ket{u}\in\complex^2$. This implies that the rank of
$\Phi(\out{u}{u})$ is independent of $\ket{u}\in\complex^2$.

Case (i). If $\rank(\Phi(\out{u}{u}))=2$, then by the spectral
decomposition theorem, it follows that
$$
\Phi(\out{u}{u}) = \mu \out{\alpha}{\alpha} + (1-\mu)
\out{\beta}{\beta},
$$
where $\mu\in(0,1)$. Moreover, there is a unitary $V$ such that
$V\ket{0}=\ket{\alpha}$ and $V\ket{1}=\ket{\beta}$. This implies
that
$$
\Phi(\out{u}{u}) = V(\mu \out{0}{0} + (1-\mu) \out{1}{1})V^\dagger.
$$
If $\mu=1-\mu$, then $\mu=\tfrac12$. That is, $\Phi(\out{u}{u}) =
\tfrac12\I$, furthermore, $\Phi(\rho)=\tfrac12\I$ for all
$\rho\in\density{\complex^2}$. If $\mu\neq\tfrac12$, define $\Psi$
as follows:
$$
\Psi(X) = \frac1{1 - 2\mu}(V^\dagger\Phi(X)V-\mu\Tr{X}\I).
$$
Hence we have that
$$
\Psi(\out{u}{u}) = \out{1}{1}.
$$
By the polarization identity for operators, we see that $\Psi$ is
rank-one preserving, it is also unitary similarity preserving, thus
there is a unitary $W$ such that
$$
\Psi(X) = W X W^\dagger\quad\text{or}\quad \Psi(X) = W X^\t
W^\dagger.
$$
Now if $\Psi(X) = W X W^\dagger$, then
$$
\Phi(X) = (1 - 2\mu)UXU^\dagger + \mu\Tr{X}\I,
$$
where $U=VW$. Letting $\lambda = 1-2\mu$ gives that
$$
\Phi(X) = \lambda UXU^\dagger + (1-\lambda)\Tr{X}\tfrac12\I
$$
for $\lambda\in[-\tfrac13,1)$ since a quantum channel is completely
positive. If $\Psi(X) = W X^\t W^\dagger$, then similarly,
$$
\Phi(X) = \lambda UX^\t U^\dagger + (1-\lambda)\Tr{X}\tfrac12\I
$$
for $\lambda\in(-1,\tfrac13]$.

Case (ii). If $\rank(\Phi(\out{u}{u}))=1$, then this corresponds to
Case (i) when $\mu=0,1$. Thus
$$
\Phi(X) = \lambda UXU^\dagger +
(1-\lambda)\Tr{X}\tfrac12\I\quad\forall \lambda\in\Br{-\tfrac13,1}
$$
or
$$
\Phi(X) = \lambda UX^\t U^\dagger +
(1-\lambda)\Tr{X}\tfrac12\I\quad\forall \lambda\in[-1,\tfrac13].
$$
The theorem is proved.
\end{proof}

Based on the above result, we come up a general \emph{conjecture}:
Let $\Phi\in\trans{\cH_d}$ be a unital quantum channel. Assume that
all eigenvalues $\lambda_i$ of a fixed quantum state $\rho_0$
satisfy that $\lambda_1>\cdots>\lambda_d>0$. Then
$\Phi(\rho)\sim\Phi(\rho_0)$ whenever $\rho\sim\rho_0$  if and only
if one of the following conditions hold.
\begin{enumerate}[(i)]
\item $\Phi(\rho) = \lambda U\rho U^\dagger + (1-\lambda)\frac1d\I$ for some $\lambda\in\Br{-\frac1{d^2-1},1}$
and unitary $U\in\unitary{\cH_d}$.
\item $\Phi(\rho) = \lambda U\rho^\t U^\dagger + (1-\lambda)\frac1d\I$ for some $\lambda\in\Br{-\frac1{d-1},\frac1{d+1}}$
and unitary $U\in\unitary{\cH_d}$.
\end{enumerate}

%=============================================================================%
\section{Concluding remarks}
%=============================================================================%

In this report, we characterized those quantum channels that preserves the majorization. We found that majorization-preserving quantum channels are equivalently described as unitary equivalence preserving quantum channels. Some open problems are still left for the future research. For example,

What can be derived for $\rS(\rho) = \rS(\sigma)$, where $\rho,\sigma\in\density{\cH}$?

If we can solve this problem, then we can use it to make an analysis of classification of quantum states according to the amount of the encoded information in quantum states. We can make further analysis of the von Neumann entropy preserving quantum channels over $\cH$: $\rS(\Phi(\rho)) = \rS(\rho)$. In particular, a special case, i.e. a unit-preserving quantum channel preserves the von Neumann entropy of a fixed quantum state, is characterized in \cite{Zhang}. Or we can still consider the characterization of quantum channel $\Phi$ such that
$$
\rS(\rho) = \rS(\sigma) \Longrightarrow \rS(\Phi(\rho)) = \rS(\Phi(\sigma)).
$$

%=============================================================================%
\subsection*{Acknowledgement}
We want to express our heartfelt thanks to Fumio Hiai
for sending the reference \cite{Hiai}
to us. This project is supported by the Research Program of Hangzhou Dianzi University (KYS075612038).

%=============================================================================%

%=============================================================================%

%=============================================================================%

\end{document}